Find sinalpha if tan(alpha/2)=1/root2
Answer by stanbon(75887)

Find sin(alpha) if tan(alpha/2) = 1/sqrt(2)
------
Since tan(alpha/2) = y/x = 1/sqrt(2),
y = 1 when x = sqrt(2)
So, r = sqrt[1^2+(sqrt(2))^2] = sqrt(1+2) = sqrt(3)
----
Then sin(alpha/2) = y/r = 1/sqrt(3)
and cos(alpha/2) = x/r = sqrt(2)/sqrt(3) = (1/3)sqrt(6)
-------------
Finally, sin(alpha) = sin(2(alpha/2)) = 2sin(alpha/2)*cos(alpha/2)
= 2
---
= (2sqrt(6)/(3*sqrt(3))
------
= (2/3)[sqrt(18)/3]
---
= (2/3)sqrt(2)
====================
Cheers,
Stan H.
====================